OpenStudy (angelab97):

How many positive 3-digit integers have a unit (ones) digit of 9? how can i solve this in the easiest way possible?

OpenStudy (anonymous):

We can use Arithmetic Sequence where... a1=109 a2=119 a3=129...and so forth up to... an=999, the last term therefore common difference d=10 what is n? an=a1+(n-1)d 999=109+10(n-1) 999-109=10n-10 900=10n n=90